1.The PCR (polymerase chain reaction) protocol that is currently used in laboratories was facilitated by the discovery of a bacterium called Thermus aquaticus in a hot spring inside Yellowstone National Park, in Wyoming. This organism contains a heat-stable form of DNA polymerase known as Taq polymerase, which continues to function even after it has been heated to 95°C.

a.Why would such a heat-stable polymerase be beneficial in PCR?

b.What would happen if it weren’t heat stable?

c.How might you choose a region of DNA for a PCR primer so as to increase the temperature necessary for primer annealing (to minimize nonspecific PCR products)?

d.A PCR reaction begins with 5 double stranded segment of DNA. Estimate the number of double-stranded copies of DNA that are present after the completion of 15 amplification cycles?

2. Based on the following replication bubble, label W, X, Y, and Z as leading or lagging strands.

53 2
Answer #1

Answer 1

  • a) Such a heat stable polymerase is beneficial in PCR because PCR involves denaturation of DNA at ~95°C, which converts double stranded DNA to single stranded DNA. During this phase the hydrogen bonds holding two DNA strands are broken.
  • b) If the enzyme were not heat stable at 95°C then during the denaturation phase, enzyme would become non functional and lose its polymerising activity. This means that we won't be able to complete the PCR cycle and no products (i.e. no replicated DNA copies) would form.
  • c) Primers should be chosen such that it's ends are rich in G/C (guanine or cytosine) residue Because there is a triple hydrogen bond between guanine and cytosine, this increases melting temperature of DNA and increases efficiency. (note: guanine and cytosine are complementary to each other).
  • d) No. Of DNA after 'n' number of cycles = 2^n
  • Thus after 15 cycles, number of ds DNA copies = (2)^15 = 32,768.
  • Since there were 5 ds DNA at beginning, total number of DNA copies=5×32768= 163840.